# FEATURE ARTICLE Visualization for Trust in Machine Learning Revisited: The State of the Field in 2023

### v1

@@ Abstract
Trust in machine learning (ML) models is a major concern in leveraging these technologies for real-world applications. Yet, ML models are being deployed in different application fields, and their role in decision-making processes is growing rapidly. Fields such as healthcare and criminal justice increasingly depend on ML models to make irreversible decisions that impact human lives. However, the black-box nature of some ML models poses a threat to their adoption. Domain experts often hesitate to rely on ML models for high-risk decision-making, as the inability to understand their inner workings fosters mistrust.
In response to the outlined challenges, researchers in academia and industry have designed several innovative solutions. For example, Google’s Explainable Artificial Intelligence (AI) Cloud and Descriptive mAchine Learning EXplanations (DALEX) package aims to improve the collaboration among domain experts to address the challenges posed by the complexity of AI. Except for all the visualization techniques analyzed in this survey article, recent frameworks set a foundation for developing techniques that facilitate users in communicating and externalizing their trust explicitly across varied ML stages and in understanding the complexities of human and AI interactions. Other works bridge the significant gap between ML outputs and human cognition by promoting a cross-disciplinary approach and building a robust model designed to examine the sender’s explanation intention and its ensuing impact on the receiver’s perception.
We base this work upon the findings of our previously published survey articles and others that have stressed the need for visual analytics (VA) to improve trust and transparency in areas such as dimensionality reduction (DR), deep learning (DL), and ML in general. After our 2020 state-of-the-art report (STAR) comprising 200 techniques, we have been collecting peer-reviewed articles describing visualization techniques for enhancing trust in ML, categorizing them based on the previously established categorization schema of 18 groups and 119 categories in total, and providing the hand-curated compilation of 542 visualization techniques in an online survey browser available at:
Although this survey article follows the same styling and overall structure as our STAR published in 2020, here, we made a new analysis of the updated dataset and discuss the most recent findings that provide us with an overview of the trust in ML visualization field as of fall 2023, including insights about the visualization community. The contributions of this article are:
@@
In this survey article, we rely on the same literature search methodology and run an identical analysis as in our 2020 STAR. We also adopt the same general definition of trust by Lee and See: “the attitude that an agent will help achieve an individual’s goals in a situation characterized by uncertainty and vulnerability” towards a more detailed, multi-level model of trust in ML that consists of five trust levels: the raw data (TL1: source reliability and transparent collection process), the processed data (TL2: uncertainty awareness, equality/data bias, comparison of structures, guidance/recommendation, and outlier detection), the algorithm/learning method (TL3: familiarity, understanding/ explanation, debugging/diagnosis, refinement/steering, comparison, knowledgeability, and fairness), the concrete model(s) for a particular task (TL4: experience, in situ comparison, performance, what-if hypotheses, model bias, and model variance), and the evaluation and the subjective users’ expectations (TL5: agreement of colleagues, visualization evaluation, metrics validation/results, and user bias). More details on methodology, trust levels, and our categorization can be found in our 2020 STAR.
@@
In the remainder of this subsection, we focus on the correlations different from our previous survey article published in 2020, but for the sake of completeness, we refer the reader to Table 4 so as to check all interesting correlations found in our dataset’s entirety.
@@
Following the same methodology as in our previous articles, we have conducted a topic analysis using latent Dirichlet allocation (LDA) based on the full texts from the PDFs. This approach assigns documents to different clusters with various weights (see supplemental material) and identifies descriptive terms for each cluster. While the topic modeling outcomes are partly influenced by the chosen parameters, they offer insights that complement those obtained from our manual analysis. Thus, the topic analysis confirms and deepens our understanding of the categorized articles.
In our approach, documents get assigned into one of ten clusters according to the topmost weight, with eight descriptive terms per cluster, as shown in Table 5. Based on a thorough review of the top terms and the content of the articles, we manually assigned the topic titles for each topic. Finally, the results are visualized as a DR projection and bar charts (see Figure 5).
@@
Compared to our 2020 STAR, Topics 2, 4, 8, 9, & 10 are entirely new due to the recent advancements in graph NNs and (vision) transformers, as well as the need for simpler tree- and rule-based systems for transparent decision-making and ML explanation. The remaining topics are similar but slightly changed compared to the 2020 counterparts. Notably, Topic 1 now refers explicitly to explaining the predictions rather than the ML’s inner workings and evaluating with user studies. Topic 3 is more prominent than before, but hyperparameter tuning is currently disregarded. Explaining individual instances and using DR projections for NN explanation also greatly surged lately (Topic 5).
@@
Most of the well-known XAI approaches that involve the use of visualization for communicating results were made by experts from other fields. The increasing popularity of these model-agnostic techniques lies in their view of a model as a mathematical function. These functions can be simplified, and interpretation methods can then assess these simpler parts’ behavior. Visualization researchers can contribute to developing VA systems for improving these methods’ interactivity and visual representations, or directly upgrading the techniques’ algorithmic components with their expertise.
@@
The ML community continually invents new NN architectures or applies existing ones in different settings, while visualization researchers are slow to adopt and work on them. For example, transformer NNs have been used in computer vision settings with promising results compared to convolutional NNs, but still, the visualization community has hardly ever experimented with helping users explore them visually. Similarly, self-supervision is another area of rapid advancement that requires the visualization community’s attention, as well as prompt engineering and model checkers for “hallucinations” of large language models (LLMs).
@@
Most VA solutions, as of recently, were arguably focusing on exploratory visual analysis. There is a trend in utilizing visual representations for confirmatory analysis and hypothesis testing, as well as causal reasoning (e.g., counterfactual reasoning, causal learning, and causal inference), which could be enhanced by visualizations. Another question is how can VA systems guide users through an exhaustive multiverse analysis for choosing the optimal model strategies that will help them understand how a complex model works and decide upon a specific action plan?
@@
Although the majority of the VA systems aim to help users understand models, quantifying the uncertainty of input and output, as well as checking their robustness and sensitivity to changes, are two major challenges because people want to deploy dynamical, well-calibrated models in real-world practical scenarios. While uncertainty quantification and sensitivity analysis are popular in other fields, visualization research is still limited in checking the inputs and outputs of models, such as with the conformal prediction method that provides statistical guarantees.
@@
Effective collaboration with domain experts from diverse fields often faces substantial communication challenges. These arise from differences in terminology, expectations, and areas of expertise, demanding huge efforts to align objectives and solutions. To mitigate the need for costly user evaluations, the exploration of AI-assisted, simulated evaluation methods is very important. The state-of-the-art benchmarking datasets should also become more rigorous since they may contain non-negligible errors. The development of rational agent benchmarks that estimate the need for and benefits of visualization can also assist XAI methods evaluation. Therefore, such benchmarks potentially minimize the dependency on extensive human participation, thereby reducing associated costs.
@@
Scaling VA systems to manage numerous instances, features, and algorithms presents significant challenges, particularly in multi-class classification scenarios. Progressive VA combined with incremental learning is one notable solution, but alternative methods for designing less complex, but more robust systems are essential in real-world settings, where covariate and label shifts are common phenomena. The concurrent use of visual channels, such as color and opacity, for other encoding tasks increases the complexity. Many users find it difficult to navigate these advanced tools, especially at initial exposure. Visualization literacy can partially alleviate this problem by educating people. Additionally, complementary to existing tutorials, storytelling can further illustrate tool functionalities, while the integration of sonification and verbalization can boost user understanding and ease of use of VA systems.
@@
Many application-specific VA systems are never deployed in the real world, making them susceptible to concept and distribution drifts. This practice highlights the need for further exploration into the out-of-domain/distribution applicability of models, emphasizing the urgency to transform specialized models into more versatile ML solutions. Ensuring the accessibility and usability of VA systems for a broader user base can be achieved through various means, including the integration of visualizations into computational notebooks and offering combined programmatic API and interactive UI solutions that elicit user feedback over long periods of time. Here, the adoption of open-source practices is vital since it not only promotes community involvement and contribution, but also addresses significant challenges related to the replication and reproducibility of ML models used in VA systems.

### v2

@@ Abstract
165mm(50mm,30mm) © 2024 IEEE. This is the author’s version of the article that has been published in IEEE Computer Graphics and Applications. The final version of this record is available at: 10.1109/MCG.2024.3360881
Trust in machine learning (ML) models is a major concern in leveraging these technologies for real-world applications.1 Yet, ML models are being deployed in different application fields, and their role in decision-making processes is growing rapidly. Fields such as healthcare and criminal justice increasingly depend on ML models to make irreversible decisions that impact human lives.2 However, the black-box nature of some ML models poses a threat to their adoption. Domain experts often hesitate to rely on ML models for high-risk decision-making, as the inability to understand their inner workings fosters mistrust.3
In response to the outlined challenges, researchers in academia and industry have designed several innovative solutions. For example, Google’s Explainable Artificial Intelligence (AI) Cloud and Descriptive mAchine Learning EXplanations (DALEX) package aims to improve the collaboration among domain experts to address the challenges posed by the complexity of AI. Except for all the visualization techniques analyzed in this survey article, recent frameworks set a foundation for developing techniques that facilitate users in communicating and externalizing their trust explicitly across varied ML stages and in understanding the complexities of human and AI interactions.4,5 Other works bridge the significant gap between ML outputs and human cognition by promoting a cross-disciplinary approach and building a robust model designed to examine the sender’s explanation intention and its ensuing impact on the receiver’s perception.6,7
We base this work upon the findings of our previously published survey articles8,9 and others that have stressed the need for visual analytics (VA) to improve trust and transparency in areas such as dimensionality reduction (DR),11 deep learning (DL),12,13 and ML in general.14,15 After our 2020 state-of-the-art report (STAR) comprising 200 techniques,9 we have been collecting peer-reviewed articles describing visualization techniques for enhancing trust in ML, categorizing them based on the previously established categorization schema of 18 groups and 119 categories in total, and providing the hand-curated compilation of 542 visualization techniques in an online survey browser available at:
Although this survey article follows the same styling and overall structure as our STAR published in 2020,9 here, we made a new analysis of the updated dataset and discuss the most recent findings that provide us with an overview of the trust in ML visualization field as of fall 2023, including insights about the visualization community. The contributions of this article are:
@@
In this survey article, we rely on the same literature search methodology and run an identical analysis as in our 2020 STAR.9 We also adopt the same general definition of trust by Lee and See:10 “the attitude that an agent will help achieve an individual’s goals in a situation characterized by uncertainty and vulnerability” towards a more detailed, multi-level model of trust in ML that consists of five trust levels: the raw data (TL1: source reliability and transparent collection process), the processed data (TL2: uncertainty awareness, equality/data bias, comparison of structures, guidance/recommendation, and outlier detection), the algorithm/learning method (TL3: familiarity, understanding/ explanation, debugging/diagnosis, refinement/steering, comparison, knowledgeability, and fairness), the concrete model(s) for a particular task (TL4: experience, in situ comparison, performance, what-if hypotheses, model bias, and model variance), and the evaluation and the subjective users’ expectations (TL5: agreement of colleagues, visualization evaluation, metrics validation/results, and user bias). More details on methodology, trust levels, and our categorization can be found in our 2020 STAR.9
@@
Domain | 542 | |
Biology | 49 | -5% |
Business | 87 | +6% |
Computer Vision | 180 | +3% |
Computers | 9 | -1% |
Health | 90 | +2% |
Humanities | 81 | -6% |
Nutrition | 16 | -1% |
Simulation | 19 | 0% - |
Social / Socioeconomic | 58 | 0% - |
Other | 230 | -5% |
Target Variable | 542 | |
Binary (categorical) | 85 | -4% |
Multi-class (categorical) | 297 | -9% |
Multi-label (categorical) | 21 | -1% |
Continuous (regression problems) | 60 | 0% - |
Other | 147 | +8% |
ML Methods | 542 | |
Convolutional Neural Network (CNN) | 72 | 0% - |
Deep Convolutional Network (DCN) | 12 | -2% |
Deep Feed Forward (DFF) | 11 | -3% |
Deep Neural Network (DNN) | 55 | 0% - |
Deep Q-Network (DQN) | 11 | -3% |
Generative Adversarial Network (GAN) | 13 | -3% |
Long Short-Term Memory (LSTM) | 26 | -2% |
Recurrent Neural Network (RNN) | 31 | -3% |
Variational Auto-Encoder (VAE) | 22 | -3% |
Other (DL methods) | 92 | +6% |
Linear (DR) | 81 | -14% |
Non-linear (DR) | 94 | -9% |
Bagging (ensemble learning) | 74 | 0% - |
Boosting (ensemble learning) | 30 | 0% - |
Stacking (ensemble learning) | 10 | -1% |
Other (generic) | 233 | -6% |
ML Types | 500 | |
Classification (supervised) | 288 | -3% |
Regression (supervised) | 55 | 0% - |
Other (supervised) | 32 | +2% |
Association (unsupervised) | 14 | 0% - |
Clustering (unsupervised) | 69 | -8% |
Dimensionality Reduction (unsupervised) | 104 | -14% |
Classification (semi-supervised) | 40 | 0% - |
Clustering (semi-supervised) | 16 | 0% - |
Classification (reinforcement) | 4 | 0% - |
Control (reinforcement) | 11 | 0% - |
Color Legend: 0 articles 271 articles 542 articles
@@
Visual Variable | 542 | |
Color | 540 | +2% |
Opacity | 294 | +12% |
Position / Orientation | 83 | -14% |
Shape | 89 | -3% |
Size | 164 | -4% |
Texture | 30 | -3% |
Evaluation | 542 | |
Standard | 132 | +5% |
Comparative | 27 | -1% |
Before / During Development | 140 | +8% |
After Development | 175 | +8% |
Not Evaluated | 196 | -15% |
Trust Levels (TL) 1–5 | 542 | |
Source Reliability | 43 | +2% |
Transparent Collection Process | 21 | +1% |
Uncertainty Awareness | 84 | 0% - |
Equality / Data Bias | 57 | +3% |
Comparison (of Structures) | 271 | +7% |
Guidance / Recommendations | 142 | +3% |
Outlier Detection | 174 | -2% |
Familiarity | 24 | +2% |
Understanding / Explanation | 314 | +10% |
Debugging / Diagnosis | 129 | -3% |
Refinement / Steering | 164 | -5% |
Comparison | 114 | -10% |
Knowledgeability | 47 | +4% |
Fairness | 29 | +2% |
Experience | 31 | +2% |
In Situ Comparison | 135 | -2% |
Performance | 338 | +8% |
What-if Hypotheses | 87 | -4% |
Model Bias | 58 | -1% |
Model Variance | 34 | -2% |
Agreement of Colleagues | 21 | -1% |
Visualization Evaluation | 301 | +12% |
Metrics Validation / Results | 413 | +11% |
User Bias | 24 | 0% - |
Target Group | 542 | |
Beginners | 104 | -2% |
Practitioners / Domain Experts | 424 | -3% |
Developers | 113 | +3% |
ML Experts | 204 | +1% |
Symbol Legend:
+/-[1 – 8]% /
+/-[9 – 16]% /
+/-[17 – 24]% /
@@
In the remainder of this subsection, we focus on the correlations different from our previous survey article published in 2020,9 but for the sake of completeness, we refer the reader to Table 4 so as to check all interesting correlations found in our dataset’s entirety.
@@
Following the same methodology as in our previous articles,8,9 we have conducted a topic analysis using latent Dirichlet allocation (LDA) based on the full texts from the PDFs. This approach assigns documents to different clusters with various weights (see supplemental material) and identifies descriptive terms for each cluster. While the topic modeling outcomes are partly influenced by the chosen parameters, they offer insights that complement those obtained from our manual analysis. Thus, the topic analysis confirms and deepens our understanding of the categorized articles.
In our approach, documents get assigned into one of ten clusters according to the topmost weight, with eight descriptive terms per cluster, as shown in Table 5. Based on a thorough review of the top terms and the content of the articles, we manually assigned the topic titles for each topic. Finally, the results are visualized as a DR projection and bar charts (see Figure 5).8,9
@@
Compared to our 2020 STAR,9 Topics 2, 4, 8, 9, & 10 are entirely new due to the recent advancements in graph NNs and (vision) transformers, as well as the need for simpler tree- and rule-based systems for transparent decision-making and ML explanation. The remaining topics are similar but slightly changed compared to the 2020 counterparts. Notably, Topic 1 now refers explicitly to explaining the predictions rather than the ML’s inner workings and evaluating with user studies. Topic 3 is more prominent than before, but hyperparameter tuning is currently disregarded. Explaining individual instances and using DR projections for NN explanation also greatly surged lately (Topic 5).
@@
Most of the well-known XAI approaches that involve the use of visualization for communicating results were made by experts from other fields. The increasing popularity of these model-agnostic techniques lies in their view of a model as a mathematical function. These functions can be simplified, and interpretation methods can then assess these simpler parts’ behavior. Visualization researchers can contribute to developing VA systems for improving these methods’ interactivity and visual representations,16 or directly upgrading the techniques’ algorithmic components with their expertise.17
@@
The ML community continually invents new NN architectures or applies existing ones in different settings, while visualization researchers are slow to adopt and work on them. For example, transformer NNs have been used in computer vision settings with promising results compared to convolutional NNs, but still, the visualization community has hardly ever experimented with helping users explore them visually.18,19 Similarly, self-supervision is another area of rapid advancement that requires the visualization community’s attention,20 as well as prompt engineering and model checkers for “hallucinations” of large language models (LLMs).21
@@
Most VA solutions, as of recently, were arguably focusing on exploratory visual analysis. There is a trend in utilizing visual representations for confirmatory analysis and hypothesis testing, as well as causal reasoning (e.g., counterfactual reasoning, causal learning, and causal inference), which could be enhanced by visualizations.22 Another question is how can VA systems guide users through an exhaustive multiverse analysis for choosing the optimal model strategies that will help them understand how a complex model works and decide upon a specific action plan?23
@@
Although the majority of the VA systems aim to help users understand models, quantifying the uncertainty of input and output, as well as checking their robustness and sensitivity to changes, are two major challenges because people want to deploy dynamical, well-calibrated models in real-world practical scenarios.24 While uncertainty quantification and sensitivity analysis are popular in other fields, visualization research is still limited in checking the inputs and outputs of models, such as with the conformal prediction method that provides statistical guarantees.25
@@
Effective collaboration with domain experts from diverse fields often faces substantial communication challenges. These arise from differences in terminology, expectations, and areas of expertise, demanding huge efforts to align objectives and solutions. To mitigate the need for costly user evaluations, the exploration of AI-assisted, simulated evaluation methods is very important. The state-of-the-art benchmarking datasets should also become more rigorous since they may contain non-negligible errors.26 The development of rational agent benchmarks that estimate the need for and benefits of visualization can also assist XAI methods evaluation.27 Therefore, such benchmarks potentially minimize the dependency on extensive human participation, thereby reducing associated costs.
@@
Scaling VA systems to manage numerous instances, features, and algorithms presents significant challenges, particularly in multi-class classification scenarios. Progressive VA combined with incremental learning is one notable solution, but alternative methods for designing less complex, but more robust systems are essential in real-world settings, where covariate and label shifts are common phenomena.28 The concurrent use of visual channels, such as color and opacity, for other encoding tasks increases the complexity. Many users find it difficult to navigate these advanced tools, especially at initial exposure. Visualization literacy can partially alleviate this problem by educating people. Additionally, complementary to existing tutorials, storytelling can further illustrate tool functionalities, while the integration of sonification and verbalization can boost user understanding and ease of use of VA systems.
@@
Many application-specific VA systems are never deployed in the real world, making them susceptible to concept and distribution drifts.29 This practice highlights the need for further exploration into the out-of-domain/distribution applicability of models, emphasizing the urgency to transform specialized models into more versatile ML solutions.30 Ensuring the accessibility and usability of VA systems for a broader user base can be achieved through various means,31 including the integration of visualizations into computational notebooks32 and offering combined programmatic API and interactive UI solutions that elicit user feedback over long periods of time.33,34 Here, the adoption of open-source practices is vital since it not only promotes community involvement and contribution, but also addresses significant challenges related to the replication and reproducibility of ML models used in VA systems.